# Beyond Trend and Periodicity: Guiding Time Series Forecasting with Textual Cues

### v1

@@ Abstract
This work introduces a novel Text-Guided Time Series Forecasting (TGTSF) task. By integrating textual cues, such as channel descriptions and dynamic news, TGTSF addresses the critical limitations of traditional methods that rely purely on historical data. To support this task, we propose TGForecaster, a robust baseline model that fuses textual cues and time series data using cross-attention mechanisms. We then present four meticulously curated benchmark datasets to validate the proposed framework, ranging from simple periodic data to complex, event-driven fluctuations. Our comprehensive evaluations demonstrate that TGForecaster consistently achieves state-of-the-art performance, highlighting the transformative potential of incorporating textual information into time series forecasting. This work not only pioneers a novel forecasting task but also establishes a new benchmark for future research, driving advancements in multimodal data integration for time series models.
@@ 1 Introduction
To achieve higher forecasting accuracy, it is essential to provide models with additional, critical information—such as holiday information, consumer sentiment, and climate changes (depending on the specific forecasting task)—that are not present in the time series data itself. Without this additional information, models cannot extract the necessary causal relationships, leading to oversimplification or overfitting. Moreover, the lack of understanding of channels and underlying systems prevents the modeling of inter-channel relationships, often resulting in unreliable forecasts.
As such valuable information is usually hard to quantify as auxiliary time series data, we consider them as textual cues and introduce a novel task for TSF: Text-Guided Time Series Forecasting (TGTSF). TGTSF leverages two main components: channel descriptions and dynamic news messages. Channel descriptions provide static knowledge about the underlying systems, enabling the model to differentiate between channels and better understand inter-channel correlations. News messages offer dynamic and external insights, helping the model adjust to shifts in data distribution caused by external events. By modeling the joint distribution of future values conditioned on these textual cues, TGTSF enhances forecasting accuracy and reliability. Furthermore, TGTSF enables scenario-based forecasting, which is particularly valuable for business decision-making. For instance, a company can forecast sales based on the assumption of a successful marketing campaign, incorporating relevant news and channel-specific descriptions to refine predictions.
We then introduce TGForecaster, a Transformer-based baseline multimodal model designed to validate the proposed TGTSF framework. Utilizing a cross-attention mechanism, TGForecaster fuses information from textual cues and time series data to guide the forecasting process. To rigorously evaluate and demonstrate the efficacy of our model, we carefully craft four benchmark datasets categorized into synthetic, captioned existing, and real-world datasets. Each dataset is uniquely designed to validate the task and test the model’s capability from different perspectives. Specifically, the synthetic dataset features simple yet variable time series patterns, such as sinusoidal waves with random frequency changes, to test the model’s ability to utilize textual cues in forecasting and mitigating distribution shifts. Two existing datasets are captioned with different levels of textual information, used to challenge the model in complex, non-periodic environments by integrating detailed external data. Lastly, our real-world dataset, collected from actual sales scenarios, showcases the practical application value of TGTSF.
@@ 2 Insights and Motivation
2.1 Time Series Forecasting Suffers from Information Insufficiency
Time series forecasting typically involves predicting future segments based on historical data, specifically forecasting subsequent segments from prior ones. However, the inherent sparsity of information within time series data makes achieving accurate forecasts challenging, or even impossible at times. Standard decomposition of time series data identifies three main components: trend, periodicity, and noise (Box et al., 2015). The noise component is intrinsically unpredictable. Trends, which are slowly changing patterns, are often impacted by external events that influence the underlying system, making it unpredictable without external information. As a result, only the periodicity component tends to be reliably predictable.
Recent advancements in linear models have validated these challenges. For instance, the DLinear (Zeng et al., 2023) model utilizes a simple linear layer to effectively capture basic periodic patterns, outperforming more complex transformer-based models. Similarly, FITS (Xu et al., 2023), designed to simply extract periodicity, achieves comparable or superior results to state-of-the-art (SOTA) methods using fewer than 10k parameters. Further research demonstrates that employing a closed-form forecasting matrix (Toner & Darlow, 2024), calculated from training data, can produce SOTA outcomes. These findings indicate that the minimal information in time series data allows models to recognize patterns with few or no learnable parameters, suggesting that training larger models with insufficient information can lead to severe overfitting.
The issue of information insufficiency extends beyond the data itself to the absence of external information. External factors like holidays, consumer sentiment, or climate changes can significantly alter patterns in time series data, affecting variables such as electricity usage, sales trends, or regional weather conditions. However, these factors are not contained within the time series data, hindering models from incorporating causal relationships. This deficiency can compromise model training, causing models to either learn overly simplified average shortcut or overfit based on the training data, as shown in Fig. 1. For example, if a dataset frequently presents days without rain, the model may consistently predict no rain for future days. However, rainfall predictions are largely influenced by external factors like climate change, not included in the time series, leading models to opt for shortcuts that minimize overall loss by predicting the most common outcome.
@@ 2.1 Time Series Forecasting Suffers from Information Insufficiency
2.2 Compromise with Information Insufficiency: RIN & Weight Sharing
Researchers have developed various methods to tackle information insufficiency in time series forecasting. One such method is weight sharing (Nie et al., 2023). This technique treats all input channels as equivalent, leveraging common patterns across all channels to alleviate the problem of insufficient training samples. While this approach enhances performance by learning from the shared periodicities typical within a single dataset, it often results in generalized outcomes. Since the model is trained to minimize loss across all channel distributions with a single set of weights,it prioritizes common patterns, failing to capture unique distributions for each channel, as illustrated in Fig. 1, which may lead to averaged results.
Figure 1: Left Panel: The traditional TSF model, when faced with distribution shifts due to external events, tends to learn an averaged shortcut, as shown in the leftmost figure. Conversely, the TGTSF task utilizes text inputs to effectively generalize and adapt to these shifts, as depicted in the second figure. Right Panel: The rightmost figure illustrates the effect of text guidance on adapting to different channel distributions. Weight Sharing (WS) method helps the model capture only common knowledge among channels, while RIN addresses distribution shifts by normalizing instances, thus learning a larger portion of the distribution. TGTSF, however, leverages text to directly learn the entire distribution, enabling the model to generalize to a specific channel distribution with corresponding text as the condition.
Another strategy is Reversible Instance Normalization (RIN) (Kim et al., 2021), which normalizes each instance to align mean and variance across the dataset, as shown in Fig. 1. RIN effectively reduces distribution shifts caused by trends or external events, aiding in the model’s ability to learn a unified distribution. However, while it standardizes data inputs, RIN also strips away essential information such as trend intensity and amplitude variations, which diminishes the model’s capacity to detect relative biases or amplitude shifts beyond the training scope.
@@ 2.2 Compromise with Information Insufficiency: RIN & Weight Sharing
2.3 Guide the Time Series Forecasting with Text
@@
To address these challenges, incorporating multimodal textual information is essential. This approach not only allows models to react to historical data but also to proactively adapt to potential future events and trends. It enhances forecasting by providing insights into the underlying dynamics and causal relationships often missing from time series data alone. Guiding time series forecasting with textual data also helps clarify inter-channel relationships, reflecting systemic correlations or domain-specific knowledge that are crucial for accurate predictions. This textual context enhances models’ ability to interpret complex data interactions, vital in environments where data channels are interconnected.
@@ 3 TGTSF Task Formulation
4 TGTSF Baseline Model
To validate the TGTSF task, we developed TGForecaster, a streamlined transformer model designed for multimodal fusion with cross-attention. Illustrated in Fig. 2, this model harnesses textual information to enhance the accuracy and relevance of time series forecasting. It employs a reimplemented PatchTST encoder for processing time series data and leverages large, pretrained text models for pre-embedding textual inputs into vector sequences across the time dimension, allowing for effective modality fusion in the embedding space.
@@ 4 TGTSF Baseline Model
A key innovation in TGForecaster is the integration of a cross-attention layer to achieve the ’text guided channel independent’. This layer is pivotal in fusing news content with channel descriptions. Taking cues from the attention mechanism’s original use in recommendation systems, we treat the news as both the key and value, with the channel descriptions serving as queries. Through this mechanism, we calculate the relevance of each news item to every channel. The resulting scores are then used to aggregate the news, culminating in a composite embedding for each channel. Another crucial aspect of our design is the modality fusion, achieved through a separate layer of cross-attention where text embeddings function as query, while temporal embeddings act as the key and value. This ensures that the output dimensions align naturally with the query which has the length of forecasting horizon.
As decoder, we use the token-wise decoding to avoid the overfitting on the last layer as observed in previous work (Lee et al., 2023). Each token is project back to its original patch length and weighted sumed to get the final forecasting result. More importantly, we do not use the RIN in TGForecaster. And our experiment shows it will worsen the performance when the information is sufficient, as shown in Appendix D.
@@ 6.1 Evaluation on Toy Dataset
Figure 3 illustrates a segment of the toy dataset where the frequency changes within the forecasting horizon. In this visualization, the PatchTST model maintains the frequency observed in the look-back window, indicating an inability to adapt to new frequencies. Similarly, DLinear displays a collapsed pattern. Notably, without news input, TGForecaster shows the same behavior of PatchTST, suggesting that it relies solely on its time series encoder in the absence of textual cues. The models without external data also highlights the decreasing amplitude of predictions, a tendency to revert to safer, average predictions when faced with uncertainty in the far future—an issue known as the ’average shortcut’ phenomenon discussed in Section 2.1. Conversely, with the integration of external text information, TGForecaster adeptly adapts to new frequencies at the appropriate moments, demonstrating the substantial benefits of incorporating textual data into the forecasting process.
@@ 6.3 Evaluation on Weather-Captioned Dataset
We follow the experiment settings in previous works as follows: forecasting horizon $H\in\{96,192,336,720\}$, look back window length of 360. We trained all the models on the Weather-Medium and -large dataset with their setting on the original weather dataset accordingly. And the TGForecaster is trained with the captioned version. We will report accordingly in Tab. 3.
@@
As demonstrated in Tab. 3, TGForecaster significantly outperforms other models across both the Weather-Medium and Weather-Large datasets, substantiating the efficacy of incorporating external text information in addressing the information insufficiency inherent in TSF models. Notably, despite its simplicity, the FITS model performs exceptionally well, second only to TGForecaster. This suggests that simply adding more training time series to complex models does not necessarily lead to improved performance, as more time series does not mean more information. The results highlight that strategic integration of textual data can provide a more substantial performance boost than merely increasing the quantity of time series data. Further, we also report the channel-wise performance in Appendix B. We notice a groundbreaking performance boost of over 60% on certain channels which were not predictable with historical time series data alone.
@@
We conducted ablation studies to evaluate the impact of different embedding models and the integration of textual inputs on the performance of TGForecaster. We tested three embedding models: OpenAI Embedding, paraphrase-MiniLM-L6, and all-mpnet-base [^3] Sentence Transformer: https://huggingface.co/sentence-transformers. The results, presented in Tab. 4, indicate minimal performance differences between the embedding models. Notably, the removal of channel descriptions led to a significant performance decrease, underscoring the model’s reliance on this feature for distinguishing between channels. Similarly, omitting news text resulted in performance dropping to levels comparable to the baseline PatchTST model, confirming that the improvements in forecasting accuracy are primarily driven by the inclusion of external textual information. This observation validates our hypothesis that textual data plays a crucial role in compensating for information deficiencies in traditional time series forecasting.
@@
The second channel, time of raining, measured in seconds per 10 minutes, presents unique challenges: it lacks periodicity, displays no causality, and often exists in a binary state. Traditional models like PatchTST often fail, typically predicting no rainfall - the averaged shortcut. TGForecaster adjusts its forecasts based on text prompts about upcoming rain. However, it sometimes misses predictions when the text data does not perfectly align with actual weather events, as noted with the missed second rainfall period.
@@
Additionally, a controllability test involving the swapping of news inputs for the second and fourth days illustrates TGForecaster’s responsive accuracy: it predicts rain on the fourth day and clear conditions on the second, aligning forecasts with the altered news data. This test underscores the model’s adaptability and the effectiveness of text-guided forecasting. See Appendix for captions.
@@ 6.4 Evaluation on Steam-100 Dataset
We evaluate the performance of the TGForecaster on our Steam-100 dataset, utilizing an input window of 60 days and an output horizon of 14 days. The findings indicate that the TGForecaster outperforms baseline models such as PatchTST, achieving a performance enhancement of over $12.6\%$. This superior performance is consistently observed across over $59.6\%$ of all games, ranking as the best among all evaluated methods. Given the significant stylistic variations among different game developers and the potential risk of temporal distribution shifts, traditional time series forecasting methods often struggle to capture the commonalities in temporal features. In contrast, the TGForecaster leverages textual information to significantly augment its predictive capabilities. We report the full result in the Appendix.

### v2

@@ Abstract
This work introduces a novel Text-Guided Time Series Forecasting (TGTSF) task. By integrating textual cues, such as channel descriptions and dynamic news, TGTSF addresses the critical limitations of traditional methods that rely purely on historical data. To support this task, we propose TGForecaster, a robust baseline model that fuses textual cues and time series data using cross-attention mechanisms. We then present four meticulously curated benchmark datasets to validate the proposed task, ranging from simple periodic data to complex, event-driven fluctuations. Our comprehensive evaluations demonstrate that TGForecaster consistently achieves state-of-the-art performance, highlighting the transformative potential of incorporating textual information into time series forecasting. This work not only pioneers a novel forecasting task but also establishes a new benchmark for future research, driving advancements in multimodal data integration for time series models.
@@ 1 Introduction
To achieve higher forecasting accuracy, it is essential to provide models with additional, critical information—such as holiday information, consumer sentiment, and climate changes (depending on the specific forecasting task)—that are not present in the time series data itself. Without this additional information, models cannot extract the necessary causal relationships, leading to oversimplification or overfitting. Moreover, the lack of insight about channels and underlying systems prevents the modeling of inter-channel relationships, often resulting in unreliable forecasts.
As such valuable information is usually hard to quantify as auxiliary time series data, we consider them as textual cues and introduce a novel task for TSF: Text-Guided Time Series Forecasting (TGTSF). TGTSF leverages two main components: channel descriptions and dynamic news messages. Channel descriptions provide static knowledge about the underlying systems, enabling the model to differentiate between channels and better understand inter-channel correlations. News messages offer dynamic and external insights, helping the model adjust to shifts in data distribution caused by external events. By modeling the joint distribution of future values conditioned on these textual cues, TGTSF has the potential to enhance forecasting accuracy and reliability. Furthermore, TGTSF enables scenario-based forecasting, which is particularly valuable for business decision-making. For instance, a company can forecast sales based on the assumption of a successful marketing campaign, incorporating relevant news and channel-specific descriptions to refine predictions.
We then introduce TGForecaster, a simple Transformer-based baseline multimodal model designed to realize the potential of the proposed TGTSF framework. Utilizing a cross-attention mechanism, TGForecaster fuses information from textual cues and time series data to guide the forecasting process. To rigorously evaluate and demonstrate the efficacy of our model, we carefully craft four benchmark datasets categorized into synthetic, captioned existing, and real-world datasets. Each dataset is uniquely designed to validate the task and test the model’s capability from different perspectives. Specifically, the synthetic dataset features simple yet variable time series patterns, such as sinusoidal waves with random frequency changes, to test the model’s ability to utilize textual cues in forecasting and mitigating distribution shifts. Two existing datasets are captioned with different levels of textual information, used to challenge the model in complex, non-periodic environments by integrating detailed external data. Lastly, our real-world dataset, collected from actual sales scenarios, showcases the practical application value of TGTSF.
@@ 2 Insights and Motivation
2.1 Existing TSF Solutions Suffer from Information Insufficiency
Time series forecasting typically involves predicting future segments based on historical data, specifically forecasting subsequent time series segments from prior ones. However, the inherent sparsity of information within time series data makes achieving accurate forecasts challenging, or even impossible at times. Standard decomposition of time series data identifies three main components: trend, periodicity, and noise (Box et al., 2015). The noise component is intrinsically unpredictable. Trends, which are slowly changing patterns, are often impacted by external events that influence the underlying system, making it less predictable without external information. As a result, the periodicity component is usually the main source of reliable prediction.
Recent advancements in linear models have validated these challenges. For instance, the DLinear (Zeng et al., 2023) model utilizes a simple linear layer to effectively capture basic periodic patterns, outperforming most complex Transformer-based models. Similarly, FITS (Xu et al., 2023), designed to simply extract periodicity, achieves comparable or superior results to state-of-the-art (SOTA) methods using fewer than $10k$ parameters. Further research demonstrates that employing a closed-form forecasting matrix (Toner & Darlow, 2024), calculated from training data, can produce SOTA outcomes. These findings indicate that the minimal information in time series data allows models to recognize patterns with few or no learnable parameters, suggesting that training larger models with insufficient information can lead to severe overfitting.
The issue of information insufficiency extends beyond the data itself to the absence of external information. External factors like holidays, consumer sentiment, or climate changes can significantly alter patterns in time series data, affecting variables such as electricity usage, sales trends, or regional weather conditions. However, these factors are not contained within the time series data, hindering models from incorporating causal relationships. This deficiency can compromise model training, causing models to either learn overly simplified average shortcut or overfit on the training data, as shown in Fig. 1. For example, if a dataset frequently presents days without rain, the model may consistently predict no rain for future days. However, rainfall predictions are largely influenced by external factors like climate change, not included in the time series, leading models to opt for shortcuts that minimize overall loss by predicting the most common outcome.
@@ 2.1 Existing TSF Solutions Suffer from Information Insufficiency
2.2 RIN & Weight Sharing are Compromises under Information Insufficiency
Researchers have developed various methods to tackle information insufficiency in time series forecasting. One such method is Reversible Instance Normalization (RIN) (Kim et al., 2021), which normalizes each instance to align mean and variance across the dataset, as shown in Fig. 1. RIN effectively reduces distribution shifts caused by trends shifting or external events, forcing the model to learn a unified distribution. While it standardizes data inputs, RIN also strips away essential information such as trend intensity and amplitude variations, which diminishes the model’s capacity to detect relative biases or amplitude shifts beyond the training scope. However, RIN merely hides the issue of information insufficiency by neutralizing the effects of external events.
Figure 1: Left Panel: The traditional TSF model, when faced with distribution shifts due to external events, tends to learn an averaged shortcut for a smaller global loss, as shown in the leftmost figure. Conversely, the TGTSF task utilizes text inputs to eliminate the uncertainty and effectively generalize to these shifts, as depicted in the second figure. Right Panel: The rightmost figure illustrates the effect of text guidance on adapting to different channel distributions. Weight Sharing (WS) method helps the model capture only common knowledge among channels, while RIN addresses distribution shifts by normalizing instances, thus learning a larger portion of the distribution. TGTSF, however, leverages text to directly learn the entire distribution, enabling the model to generalize to a specific channel distribution with corresponding text as the condition.
Another strategy is weight sharing (Nie et al., 2023; Zeng et al., 2023; Xu et al., 2023). This technique treats all input channels as equivalent, leveraging common patterns across all channels to alleviate the problem of limited training samples. In other words, this approach enhances performance by learning from the shared periodicities across all the channels, typical within a single dataset. However, since the model is trained to minimize loss across all channel distributions with a single set of weights, it prioritizes common patterns, failing to capture unique distributions for each channel, as illustrated in Fig. 1, which may lead to averaged results. These average results further reveal that the model lacks sufficient information to distinguish between different channels effectively.
@@ 2.2 RIN & Weight Sharing are Compromises under Information Insufficiency
2.3 Motivation for The Proposed TGTSF Framework
@@
To address these challenges, it is essential to include text as an additional input modality. This approach not only allows models to react to historical data but also to proactively adapt to potential future events and trends. It enhances forecasting by providing insights into the underlying dynamics and causal relationships often missing from time series data alone. Guiding time series forecasting with textual data also helps clarify inter-channel relationships, reflecting systemic correlations or domain-specific knowledge that are crucial for accurate predictions. This textual context enhances models’ ability to interpret complex data interactions, vital in environments where data channels are interconnected.
@@ 3 TGTSF Task Formulation
4 TGForecaster: A Baseline Model for The TGTSF Task
To validate the TGTSF task, we developed TGForecaster, a streamlined transformer model designed for multimodal fusion with cross-attention. Illustrated in Fig. 2, this model harnesses textual information to enhance the accuracy and relevance of time series forecasting. It employs a reimplemented PatchTST encoder for processing time series data and leverages off-the-shelf, pretrained text models for pre-embedding textual inputs into vector sequences across the time dimension, allowing for effective modality fusion in the embedding space.
@@ 4 TGForecaster: A Baseline Model for The TGTSF Task
A key innovation in TGForecaster is the integration of a cross-attention layer to achieve the ’text guided channel independent’. This layer is pivotal in fusing news content with channel descriptions. Taking cues from the attention mechanism’s original use in recommendation systems, we treat the news as both the key and value, with the channel descriptions serving as queries. Through this mechanism, we calculate the relevance of each news item to every channel. The resulting scores are then used to aggregate the news, culminating in a composite embedding for each channel. Another crucial aspect of our design is the modality fusion, achieved through a separate layer of cross-attention where text embeddings function as query, while temporal embeddings act as the key and value. This ensures that the output dimensions align naturally with the query which has the length of forecasting horizon. The attention maps for these two cross-attention layers are provided in Appendix E. These maps compellingly demonstrate that the layers perform in accordance with our design specifications.
As decoder, we use the token-wise decoding to avoid the overfitting on the last layer as observed in previous work (Lee et al., 2023). Each token is project back to its original patch length and weighted sumed to get the final forecasting result. Notably, TGForecaster does not incorporate the RIN. Our experiments, detailed in Appendix D, demonstrate that the inclusion of RIN degrades performance when sufficient information is already available.
@@ 6.1 Evaluation on Toy Dataset
Figure 3 illustrates a segment of the toy dataset where the frequency changes within the forecasting horizon. In this visualization, the PatchTST model maintains the frequency observed in the look-back window, indicating an inability to adapt to new frequencies. Similarly, DLinear displays a collapsed pattern. Notably, without news input, TGForecaster shows the same behavior of PatchTST, suggesting that it relies solely on its time series encoder in the absence of textual cues. The models without external data also highlights the decreasing amplitude of predictions, a tendency to revert to safer, average predictions when faced with uncertainty in the far future—an issue known as the ’average shortcut’ phenomenon discussed in Section 2.1. Conversely, with the integration of external text information eliminating uncertainty, TGForecaster adeptly adapts to new frequencies at the appropriate moments, demonstrating the substantial benefits of incorporating textual data into the forecasting process.
@@ 6.3 Evaluation on Weather-Captioned Dataset
We follow the experiment settings in previous works as follows: forecasting horizon $H\in\{96,192,336,720\}$, look back window length of 360. We trained all other models on the Weather-Medium and -large dataset with their setting on the original weather dataset accordingly. And the TGForecaster is trained with the captioned version. We will report accordingly in Tab. 3.
@@
As demonstrated in Tab. 3, TGForecaster significantly outperforms other models across both the Weather-Medium and Weather-Large datasets, substantiating the efficacy of incorporating external text information in addressing the information insufficiency inherent in TSF models. Notably, despite its simplicity, the FITS model performs exceptionally well, second only to TGForecaster. This suggests that simply adding more training time series to complex models does not necessarily lead to improved performance, as more time series does not necessarily contain more information. The results highlight that strategic integration of textual data can provide a more substantial performance boost than merely increasing the quantity of time series data. Further, we also report the channel-wise performance in Appendix B. We notice a groundbreaking performance boost of over 60% on certain channels which were not predictable with historical time series data alone.
@@
We conducted ablation studies to evaluate the impact of different embedding models and the integration of textual inputs on the performance of TGForecaster. We tested three embedding models: OpenAI Embedding (ope, ), paraphrase-MiniLM-L6 (Wang et al., 2020), and all-mpnet-base (Song et al., 2020) [^3] Sentence Transformer: https://huggingface.co/sentence-transformers. The results, presented in Tab. 4, indicate minimal performance differences between the embedding models. Notably, the removal of channel descriptions led to a significant performance decrease, underscoring the model’s reliance on this feature for distinguishing between channels. Similarly, omitting news text resulted in performance dropping to levels comparable to the baseline PatchTST model, confirming that the improvements in forecasting accuracy are primarily driven by the inclusion of external textual information. This observation validates our hypothesis that textual data plays a crucial role in compensating for information deficiencies in traditional time series forecasting.
@@
The second channel, time of raining, measured in seconds per 10 minutes, presents unique challenges: it lacks periodicity, displays no causality, and often exists in a binary state. Traditional models like PatchTST often fail, typically predicting no rainfall - the averaged shortcut. TGForecaster adjusts its forecasts based on text prompts about upcoming rain. However, it sometimes misses predictions when the text weather forecasting report does not perfectly align with actual weather events, as noted with the missed second rainfall period.
@@
Additionally, a controllability test involving the swapping of news inputs for the second and fourth days illustrates TGForecaster’s responsive accuracy: it predicts rain on the fourth day and clear conditions on the second, aligning forecasts with the altered news data. This test underscores the model’s adaptability and the effectiveness of text-guided forecasting.
@@ 6.4 Evaluation on Steam-100 Dataset
We evaluate the performance of the TGForecaster on our Steam-100 dataset, utilizing an input window of 60 days and an output horizon of 14 days. The findings indicate that the TGForecaster outperforms baseline models such as PatchTST, achieving a performance enhancement of over $12.6\%$. This superior performance is consistently observed across over $59.6\%$ of all games, ranking as the best among all evaluated methods. Given the significant stylistic variations among different game developers and the potential risk of temporal distribution shifts, traditional time series forecasting methods often struggle to capture the commonalities in temporal features. In contrast, the TGForecaster leverages textual information to significantly augment its predictive capabilities. We report the full result in the Appendix G.